OpenStudy (anonymous):

Which expression describes the rule of this sequence? 20, 40, 60, 80, 100........

OpenStudy (trantom):

x+20 since you add 20 each time

OpenStudy (rina.r):

The sequence is 20 so yeah x+20

OpenStudy (anonymous):

It is multiple choice 20 + n 20n 20^n n + 20 + 20

OpenStudy (anonymous):

The difference is d=20 ,so we add each number 20

OpenStudy (trantom):

A, 20 + n

OpenStudy (anonymous):

An=20+(n-1)(20)=20-20+20n=20n

OpenStudy (anonymous):

How is it 20+n if n is the term of the sequence?

OpenStudy (anonymous):

Formula An=A1+(n-1)(d)

OpenStudy (anonymous):

Thanks guys ^_^

OpenStudy (anonymous):

np